OpenStudy (anonymous):

What is the slope-intercept form of an equation of a line perpendicular to y = 2x - 3 and passing through the point (5,-1)? Question 13 options: y = 2x - 11 y= -1/2x -3/2 y= -1/2x +3/2 y = 2x + 11

OpenStudy (anonymous):

Write the equation in slope-intercept form: -5x-8y=2

OpenStudy (anonymous):

??

OpenStudy (anonymous):

Which of the following equations represents a line perpendicular to y= -4x + 1? Question 15 options: -8y = -4x -7 -8y + 2x = 1 y= 32x -8 -8y - 2x = 1

OpenStudy (anonymous):

Which of the following equations represents a line perpendicular to -6x + 7y = -1 ? Question 1 options: 30x + 35y = -1 -35x - 30y = -1 -6x - 35y = -1 7x - 6y = -1

OpenStudy (anonymous):

??